The sum of the digits of a two-digit number is 10. If the digits are reversed,the new number is one less than twice the original number. What is the number?
Answer by stanbon(48558)

The sum of the digits of a two-digit number is 10. If the digits are reversed,the new number is one less than twice the original number. What is the number?
--------
Comment: Any two-digit number can be expressed as 10t+u where
t is the ten's digit and u is the unit's digit.
----------------------
EQUATION;
t+u = 10
10u+t = 2(10t+u)-1
Simplify the 2nd equation as follows:
10u+t = 20t+2u-1
19t-8u=1
-----------
Rewrite as follows:
t+u=10
19t-8u=1
----------
t=10-u
Substitute into the 2nd equation to solve for u, as follows:
19(10-u)-8u=1
190-27u=1
27u=189
u=7
Then t=10-u=3
----------
The original number is 37
==========
Cheers,
Stan H.
: